# Universal Ahlfors–David regularity of Steiner trees

### v1

@@ Abstract
$\frac{\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x))}{\varepsilon}\leq\left(\frac{64d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{64d}{1-\rho}\right)^{d-1}$
@@ 1 Introduction
The Steiner minimal tree problem (hereafter simply the Steiner problem) is an umbrella term for an entire class of problems involving the minimization of the length of a set that connects a collection of points. We will study the Steiner problem in two specific formulations (Problems 1 and 2).
@@
It is known that the set $\mathcal{S}t$ exists (but is not always unique) and is a union of a finite set of segments. Thus, $\mathcal{S}t$ can be identified with an embedded graph whose vertex set contains $\mathcal{A}$ and whose edges are straight line segments. This graph $G$ is connected and contains no cycles, which is why the set $\mathcal{S}t$ is a tree, explaining its name. It is known that the vertices of $G$ cannot have degree greater than 3. Furthermore, only vertices corresponding to some $a_{i}$ can have degree 1 or 2; the remaining vertices have degree 3 and are called branching points (note that a point from $\mathcal{A}$ can also have degree 3). The number of branching points does not exceed n2n-2. The pairwise angles between edges meeting at any vertex are at least $2\pi/3$. Thus, three edges meet at a branching point with pairwise angles of $2\pi/3$. In any ambient dimension $d$, the three edges adjacent to a branching point lie in a two-dimensional plane. Branching points not in $\mathcal{A}$ are called Steiner points. If all vertices $a_{i}$ have degree 1, then the number of Steiner points is n2n-2 (and the converse also holds), and $\mathcal{S}t$ is called a full Steiner tree for $\mathcal{A}$. The proof of these properties of Steiner trees and other information about them can be found in the book [7] and in the work [5].
We call a local Steiner tree for a finite set $\mathcal{A}$ a connected, compact, acyclic set $S$ containing $\mathcal{A}$, such that for every point $x\in S$, there exists a neighborhood $U\ni x$ such that $S\cap U$ coincides with a Steiner tree connecting the points $(S\cap\partial U)\cup(\mathcal{A}\cap U)$. A local Steiner tree possesses the following properties of a Steiner tree: it consists of a finite set of segments, and the angles between the segments meeting at any point of $S$ are at least $2\pi/3$. We will use the fact that a connected closed subset of a local Steiner tree is also a local Steiner tree.
@@ 1.1 Notation
For a given set $X\subset\mathbb{R}^{d}$, we denote by $\overline{X}$ its closure, by $\operatorname{Int}(X)$ its interior, and by $\partial X$ its topological boundary. Let $B_{\rho}(x)$ stand for the open ball of radius $\rho$ centered at a point $x$. For given points $b$, $c\in\mathbb{R}^{d}$, we use the notation $[bc]$ for the corresponding closed line segment. The distance between $b$ and $c$ is denoted by $\operatorname{dist}(b,c)$, except in the case $d=2$ and a complex coordinate system in which we write |bc||b-c|.
@@ 1.2 General Steiner problem
If $S$ is a subset of a topological space, we say that $S$ is connected if it is not the disjoint union of two non-empty closed sets. We say that $S$ is path-connected if given any two points of $S$ there is a continuous curve joining them in $S$. We say that $S$ contains no loops if no subset of $S$ is homeomorphic to the circle $\mathbb{S}^{1}$. We say that $S$ is a topological tree if $S$ is connected and contains no loops.
Let $\mathcal{A}\subset\mathbb{R}^{d}$ and $\mathcal{S}t\subset\mathbb{R}^{d}$. We say that $\mathcal{S}t$ is a Steiner set for $\mathcal{A}$ if $\mathcal{S}t\cup\mathcal{A}$ is connected and $\mathcal{H}^{1}(\mathcal{S}t)\leq\mathcal{H}^{1}(S)$ whenever $S\subset\mathbb{R}^{d}$ and $S\cup\mathcal{A}$ is connected. We say that $\mathcal{S}t$ is a Steiner tree if $\mathcal{S}t$ is a Steiner set and is itself connected.
@@
Let $\mathcal{S}t$ be a Steiner tree of finite length for $\mathcal{A}\subset\mathbb{R}^{d}$, and assume that the ball $B_{r}(x)$ does not intersect $\mathcal{A}$. Then for any $\varepsilon>0$, the set
@@
$\mathcal{S}t^{\prime}:=\mathcal{S}t\setminus B_{\varepsilon^{\prime}}(\mathcal{A})$
consists of a finite number of segments. By choosing $\varepsilon^{\prime}<\varepsilon$ such that this finiteness holds, we obtain that $\mathcal{S}t\cap\overline{B_{r-\varepsilon}(x)}\subset\mathcal{S}t^{\prime}$. Since the ball $\overline{B_{r-\varepsilon}(x)}$ is convex, it intersects each of the segments of $\mathcal{S}t^{\prime}$ in a segment. ∎
@@ 1.3 Contribution
Fix $d>2$. Let $\mathcal{S}t$ be a Steiner tree for $\mathcal{A}$ with $\mathcal{H}^{1}(\mathcal{S}t)<\infty$. Assume that for some $s>0$ and some $x\in\mathbb{R}^{d}$, the open ball $B_{s}(x)$ has empty intersection with $\mathcal{A}$, and choose $\rho\in(0,1)$. Then
$\frac{\mathcal{H}^{1}(\mathcal{S}t\cap B_{\rho s}(x))}{s}\leq\left(\frac{64d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{64d}{1-\rho}\right)^{d-1}$
@@
It is a more delicate question whether the condition $\rho\in(0,1)$ is unavoidable here. However, Theorem 5 shows that even in the plane the statement of Corollary 2 fails for $\rho=1$.
@@
$\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x))<\varepsilon\left(\frac{64d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{64d}{1-\rho}\right)^{d-1}$
@@
In Subsection 1.4 we introduce the notion of Ahlfors–David regularity and explain why the assumption $\mathcal{H}^{1}(\mathcal{S}t)<+\infty$ and a restriction to a smaller ball are unavoidable. Section 2 contains a few classical results we use; despite the fact that some of them seem to be well-known, they are difficult to find in the required form, so we accompany them with proofs. In Section 3 we present a trickier example, showing that even in the plane we need a smaller ball to ensure that the combinatorial structure inside the ball is finite; in Subsection 4.3 we prove a certain extremal property of this example. Subsection 4.2 contains the proofs of the main results (Theorem 2 and Corollaries 2 and 3). Finally, Section 5 collects open questions.
@@ 1.4 Ahlfors–David Regularity
Thus, it is necessary to require $\mathcal{H}^{1}(\mathcal{S}t)<+\infty$. However, even in this case, the Steiner tree can be as irregular as $\mathcal{A}$; in particular, it may have an arbitrarily long length in a given ball. An alternative explanation of this phenomenon is the fact that the length of a Steiner tree for $N$ random points in the unit ball grows proportionally to $N^{\frac{d-1}{d}}$ (the maximum length has the same order of growth).
@@
It is clear that in any Euclidean space $\mathbb{R}^{d}$, we can take $c=2$, and this estimate is sharp for points in the interior of any segment from any Steiner tree and small enough $r$. In the plane, regularity with constant $C=2\pi$ is obvious (a competitor can be produced by a replacement of a part of a Steiner set inside a circle with the boundary of the circle), but finding the smallest possible value of $C$ is a non-trivial problem.
In higher dimensions it is impossible to have universal regularity in the sense of Definition 2. Indeed, if we take $N$ random points on the sphere $\mathbb{S}^{d-1}=\partial B_{1}(o)$, then the length of a Steiner tree for them is at most $N$ (and so finite). However, the sum of the radii of disjoint neighborhoods as in the analysis of Example 1 will tend to infinity as $N\to\infty$. So a constant $C$ in Definition 2 depends on $\mathcal{S}t$. On the other hand, if we consider a smaller ball $B_{\rho r}(x)$, then most of the length lies outside $B_{\rho r}(x)$ and a universal estimate like the one in Theorem 2 becomes possible.
@@ 2.1 Convex hull property
A Steiner set $\mathcal{S}t$ for a compact $\mathcal{A}\subset\mathbb{R}^{d}$ is a subset of $\operatorname{conv}\mathcal{A}$.
@@ 2.2 Combinatorics of finite locally minimal trees
The topology is called degenerate if a point from $\mathcal{A}$ has degree 3 in it. The name is due to the fact that a random tree on $n$ vertices has a degenerate topology with probability 0, and any other topology occurs with positive probability.
Note that if all vertices lie on a sphere, no locally minimal tree can have a degenerate topology.
@@
It follows from Proposition 1 that $S_{T}(\mathcal{A})$ is well-defined. Moreover, one can construct $S_{T}(\mathcal{A})$ (or show that it is impossible) in linear time $O(n)$, where $n$ is the number of terminals [9]. However, we rarely know a priori which realizable topologies correspond to Steiner trees. Although the number of possible topologies for an $n$-point configuration is finite, checking all of them can take a long time, as this number of topologies grows factorially with $n$, [5]. Furthermore, the Steiner problem is NP-hard, [4].
For a full topology $T$, we define $D(T)$ as the set of topologies that can be obtained from $T$ by contracting some edges (possibly none, i.e., $T\in D(T)$) connecting a terminal to a Steiner point (these edges must have pairwise distinct ends).
@@
Let $T$ be a full topology and $\mathcal{A}$ be a finite subset of the plane. Consider the function
@@
which is the length of the tree on the set of vertices $\mathcal{A}\cup\{y_{1},\dots,y_{n-2}\}$ with branching points $y_{1},\dots,y_{n-2}$ by straight edges and with topology $T$ (we allow $y_{i}$ to coincide with each other and with terminals, as well as edge intersections). Then $L$ has a unique local minimum, and consequently, there is at most one locally minimal tree for $\mathcal{A}$ with a topology from $D(T)$.
It is easy to see that the sets $D(T)$ form a partition of the non-degenerate topologies. Indeed, consider a non-full topology $R$ that is not degenerate. For each vertex $x$ of degree two in $R$, we add a new branching point, connect it to $x$’s neighbors (say, $y$ and $z$), and also to $x$ itself, and remove the edges $xy$, xzxz. The resulting full abstract tree $T$ is the unique one for which $R\in D(T)$.
@@ 2.3 Geometry of planar finite locally minimal trees
Let $S$ be a full locally minimal tree for a finite $\mathcal{A}\subset\mathbb{R}^{2}$. Let $q$ be a branching point adjacent to two pendant vertices of $S$, say $p_{1}$ and $p_{2}$, then the length $\mathcal{H}^{1}(S)$ is equal to the length of the connected set $S^{\prime}$ obtained from $S$ by removing the points $p_{1}$ and $p_{2}$ and the edges $p_{1}q]$ and $p_{2}q]$ and prolonging the remaining edge arriving in $q$ up to a new vertex $p$ such that $pp_{1}p_{2}$ is an equilateral triangle and (among the two possibilities) $p$ and $q$ are on opposite sides of the line $(p_{1},p_{2})$.
@@
Clearly, $\angle p_{1}pp_{2}=\pi/3$, $\angle p_{1}qp_{2}=2\pi/3$ and the line $(p_{1},p_{2})$ separates $p$ and $q$. Hence $p,p_{1},p_{2}$ and $q$ are cocircular. Then $\angle p_{1}qp=\angle p_{1}p_{2}p=\pi/3$, so $[pq]$ is a continuation of the remaining edge adjacent to $q$ in $S$. By the Ptolemy’s theorem
@@ 2.4 Stability of finite solutions
Let $\mathcal{A}=\{a_{1},\dots,a_{n}\}$, and let the shifted vertices form the set $\mathcal{B}=\{b_{1},\dots,b_{n}\}$.
Add segments $a_{i}b_{i}]$ to $\mathcal{S}t$, where $b_{i}$ is the shifted vertex; then replace all edge-segments $a_{i}x]$ with $b_{i}x]$, and consider the unique local minimum $\mathcal{S}t_{new}$ from Proposition 1. Each terminal has degree 1 or 2, so by the triangle inequality
@@
Consider an arbitrary locally minimal tree $S$ for $\mathcal{B}$ with a topology not in $D(T)$. Similarly to the previous paragraph, add segments $a_{i}b_{i}]$ to $S$, replace all edge-segments $b_{i}x]$ with $a_{i}x]$ and consider the unique local minimum $S_{old}$ from Proposition 1. Similarly, each terminal has degree 1, 2 or 3, so
@@
By assumption $\mathcal{H}^{1}(\mathcal{S}t)+\varepsilon<\mathcal{H}^{1}(S_{old})$, therefore $\mathcal{H}^{1}(\mathcal{S}t_{new})<\mathcal{H}^{1}(S)$.

### v2

@@ Abstract
$\frac{\mathcal{H}^{1}\left(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x)\right)}{\varepsilon}\leq\left(\frac{144d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{144d}{1-\rho}\right)^{d-1}$
@@ 1 Introduction
The Steiner minimal tree problem (hereafter simply the Steiner problem) is an umbrella term for an entire class of problems involving the minimization of the length of a set that connects a given collection of points. We will study the Steiner problem in two specific formulations (Problems 1 and 2).
@@
It is known that the set $\mathcal{S}t$ exists (but is not always unique) and is a union of a finite set of segments. Thus, $\mathcal{S}t$ can be identified with an embedded graph whose vertex set contains $\mathcal{A}$ and whose edges are straight line segments. This graph $G$ is connected and contains no cycles, which is why the set $\mathcal{S}t$ is a tree, explaining its name. It is known that the vertices of $G$ cannot have degree greater than 3. Furthermore, only vertices corresponding to some $a_{i}$ can have degree 1 or 2; the remaining vertices have degree 3 and are called branching points (note that a point from $\mathcal{A}$ can also have degree 3). The number of branching points does not exceed n2n-2. The pairwise angles between edges meeting at any vertex are at least $2\pi/3$. Thus, three edges meet at a branching point with pairwise angles of $2\pi/3$. In any ambient dimension $d$, the three edges adjacent to a branching point lie in a two-dimensional plane. Branching points outside $\mathcal{A}$ are called Steiner points. If all vertices $a_{i}$ have degree 1, then the number of Steiner points is n2n-2 (and the converse also holds), and $\mathcal{S}t$ is called a full Steiner tree for $\mathcal{A}$. The proof of these properties of Steiner trees and other information about them can be found in the book [10] and in the work [7].
We call a locally minimal tree for a finite set $\mathcal{A}$ a connected, compact, acyclic set $S$ containing $\mathcal{A}$, such that for every point $x\in S$, there exists a neighborhood $U\ni x$ such that $S\cap U$ coincides with a Steiner tree connecting the points $(S\cap\partial U)\cup(\mathcal{A}\cap U)$. A locally minimal tree possesses the following properties of a Steiner tree: it consists of a finite set of segments, and the angles between the segments meeting at any point of $S$ are at least $2\pi/3$.
@@ 1.1 Notation
For a given set $X\subset\mathbb{R}^{d}$, we denote by $\overline{X}$ its closure, by $\operatorname{Int}(X)$ its interior, and by $\partial X$ its topological boundary. The size of a finite set $X$ is denoted by #X\#X. Let $B_{\rho}(x)$ stand for the open ball of radius $\rho$ centered at a point $x$.
For given points $b$, $c\in\mathbb{R}^{d}$, we use the notation $[bc]$, $[bc)$ and $(bc)$ for the corresponding closed line segment, ray, and line respectively. The distance between points $b$ and $c$ is denoted by $\operatorname{dist}(b,c)$, except in the case $d=2$ and a complex coordinate system in which we write |bc||b-c|. Also, we denote the origin by $o$.
@@ 1.2 General Steiner problem
If $S$ is a subset of a topological space, we say that $S$ is connected if it is not the disjoint union of two nonempty closed sets. We say that $S$ is path-connected if for any given two points of $S$ there is a continuous curve joining them in $S$. We say that $S$ contains no loops if no subset of $S$ is homeomorphic to the circle $\mathbb{S}^{1}$. We say that $S$ is a topological tree if $S$ is connected and contains no loops.
Let $\mathcal{A}\subset\mathbb{R}^{d}$ and $\mathcal{S}t\subset\mathbb{R}^{d}$. We say that a closed set $\mathcal{S}t$ is Steiner set for $\mathcal{A}$ if $\mathcal{S}t\cup\mathcal{A}$ is connected and $\mathcal{H}^{1}(\mathcal{S}t)\leq\mathcal{H}^{1}(S)$ whenever $S\subset\mathbb{R}^{d}$ and $S\cup\mathcal{A}$ is connected. We say that $\mathcal{S}t$ is a Steiner tree if $\mathcal{S}t$ is a Steiner set and is itself connected. A Steiner tree $\mathcal{S}t$ is called full if $\mathcal{S}t\setminus\mathcal{A}$ is connected.
@@
Let $\mathcal{S}t$ be a Steiner tree of finite length for a compact set $\mathcal{A}\subset\mathbb{R}^{d}$, and assume that the ball $B_{r}(x)$ does not intersect $\mathcal{A}$. Then for any $\varepsilon>0$, the set
@@
$\mathcal{S}t_{\varepsilon^{\prime}}:=\mathcal{S}t\setminus B_{\varepsilon^{\prime}}(\mathcal{A})$
consists of a finite number of segments. By choosing $\varepsilon^{\prime}<\varepsilon$ such that this finiteness holds, we obtain that $\mathcal{S}t\cap\overline{B_{r-\varepsilon}(x)}\subset\mathcal{S}t_{\varepsilon^{\prime}}$. Since the ball $\overline{B_{r-\varepsilon}(x)}$ is convex, it intersects each of the segments of $\mathcal{S}t_{\varepsilon^{\prime}}$ in at most one segment. ∎
@@ 1.3 Contribution
Fix $d>2$. Let $\mathcal{S}t$ be a Steiner tree for $\mathcal{A}$ with $\mathcal{H}^{1}(\mathcal{S}t)<\infty$. Assume that for some $s>0$ and some $x\in\mathbb{R}^{d}$, the open ball $B_{s}(x)$ has empty intersection with $\mathcal{A}$, and choose any $\rho\in(0,1)$. Then
$\frac{\mathcal{H}^{1}(\mathcal{S}t\cap B_{\rho s}(x))}{s}\leq\left(\frac{144d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{144d}{1-\rho}\right)^{d-1}$
@@
The reasons for requiring the condition $\rho\in(0,1)$ are more delicate than those for the finiteness of $\mathcal{H}^{1}(\mathcal{S}t)$. However, Theorem 5 shows that even in the plane the statement of Corollary 2 fails for $\rho=1$.
@@
$\mathcal{H}^{1}(\mathcal{S}t_{\varepsilon}\cap B_{\rho\varepsilon}(x))<\varepsilon\left(\frac{144d}{1-\rho}\right)^{d-2}.$
@@
$\left(\frac{144d}{1-\rho}\right)^{d-1}$
@@
In Subsection 1.4 we introduce the notion of Ahlfors–David regularity and explain why the assumption $\mathcal{H}^{1}(\mathcal{S}t)<+\infty$ and the restriction to a smaller ball are unavoidable. Section 2 contains a few classical results we use; despite the fact that some of them seem to be well-known, they are difficult to find in the required form, so we accompany them with proofs. In Section 3 we present a trickier example, showing that even in the plane we need a smaller ball to ensure that the combinatorial structure inside the ball is finite; in Subsection 4.3 we prove a certain extremal property of this example. Subsection 4.2 contains the proofs of the main results (Theorem 2 and Corollaries 2 and 3). Finally, Section 5 collects open questions.
@@ 1.4 Ahlfors–David Regularity
Thus, it is necessary to require $\mathcal{H}^{1}(\mathcal{S}t)<+\infty$. However, even in this case, the Steiner tree can be as irregular as $\mathcal{A}$; in particular, it may have an arbitrarily long length in a given ball. An alternative explanation of this phenomenon is the fact that the length of a Steiner tree for $N$ random points in the unit ball in $\mathbb{R}^{d}$ grows proportionally to $N^{\frac{d-1}{d}}$ (the maximum length has the same order of growth).
@@
It is clear that in any Euclidean space $\mathbb{R}^{d}$, we can take $c=2$, and this estimate is sharp for points in the interior of any segment from any Steiner tree and small enough $r$. In the plane, regularity with constant $C=2\pi$ is obvious (a competitor can be produced by a replacement of a part of a Steiner set inside a circle with the boundary of the circle), but finding the smallest possible value of $C$ is a nontrivial problem.
In higher dimensions it is impossible to have universal regularity in the sense of Definition 2. Indeed, if we take $N$ random points on the sphere $\mathbb{S}^{d-1}=\partial B_{1}(o)$, then the length of a Steiner tree for them is at most $N$ (and so finite). However, the sum of the radii of disjoint neighborhoods as in the analysis of Example 1 will tend to infinity as $N\to\infty$. So a constant $C$ in Definition 2 depends on $\mathcal{S}t$. On the other hand, if we consider a smaller ball $B_{\rho}(o)$, then most of the length lies outside $B_{\rho}(o)$ and a universal estimate like the one in Theorem 2 becomes possible.
@@ 2.1 Convex hull property
A Steiner set $\mathcal{S}t$ for a compact $\mathcal{A}\subset\mathbb{R}^{d}$ is a subset of $\operatorname{conv}\mathcal{A}$, provided that $\mathcal{H}^{1}(\mathcal{S}t)<\infty$.
@@ 2.2 Combinatorics of finite locally minimal trees
In this subsection we consider the set of topologies with $n$ terminal points.
The topology is called degenerate if a point from $\mathcal{A}$ has degree 3 in it. The name is due to the fact that a random tree on $n$ vertices has a degenerate topology with probability 0, and any other topology occurs with positive probability. Note that if all vertices lie on a sphere, no locally minimal tree can have a degenerate topology.
@@
It follows from Proposition 1 that $S_{T}(\mathcal{A})$ is well-defined. Moreover, one can construct $S_{T}(\mathcal{A})$ (or show that it is impossible) in linear time $O(n)$, where $n$ is the number of terminals, see [9]. However, we rarely know a priori which realizable topologies correspond to Steiner trees. Although the number of possible topologies for an $n$-point configuration is finite, checking all of them can take a long time, as this number of topologies grows factorially with $n$, see [7]. Furthermore, the Steiner problem is NP-hard, see [6].
For a full topology $T$ with $n$ terminals, we define $D(T)$ as the set of topologies with $n$ terminals that can be obtained from $T$ by contracting some edges (possibly none, i.e., $T\in D(T)$) connecting a terminal to a Steiner point (these edges must have pairwise distinct ends).
@@
Let $T$ be a full topology with $n$ terminals and $\mathcal{A}$ be a finite subset of cardinality $n$ in the plane. Consider the function
@@
which is the length of the tree on the set of vertices $\mathcal{A}\cup\{y_{1},\dots,y_{n-2}\}$ with branching points $y_{1},\dots,y_{n-2}$ connected by straight edges according to the topology $T$ (we allow $y_{i}$ to coincide with each other and with terminals, as well as edge intersections). Then $L$ has a unique local minimum, and consequently, there is at most one locally minimal tree for $\mathcal{A}$ with a topology from $D(T)$.
It is easy to see that the sets $D(T)$ form a partition of the set of all non-degenerate topologies with $n$ terminals. Indeed, consider a non-full topology $R$ that is not degenerate. For each vertex $x$ of degree two in $R$, we add a new branching point, connect it to $x$’s neighbors (say, $y$ and $z$), and also to $x$ itself, and remove the edges $xy$, xzxz. The resulting full abstract tree $T$ is the unique one for which $R\in D(T)$.
@@ 2.3 Geometry of planar finite locally minimal trees
Let $S$ be a full locally minimal tree for a finite $\mathcal{A}\subset\mathbb{R}^{2}$. Let $q$ be a branching point adjacent to two pendant vertices of $S$, say $p_{1}$ and $p_{2}$. Then the length $\mathcal{H}^{1}(S)$ is equal to the length of the connected set $S^{\prime}$ obtained from $S$ by removing the points $p_{1}$ and $p_{2}$ and the edges $p_{1}q]$ and $p_{2}q]$ and prolonging the remaining edge arriving in $q$ up to a new vertex $p$ such that $pp_{1}p_{2}$ is an equilateral triangle and (among the two possibilities) $p$ and $q$ are on opposite sides of the line $(p_{1},p_{2})$.
@@
Clearly, $\angle p_{1}pp_{2}=\pi/3$, $\angle p_{1}qp_{2}=2\pi/3$ and the line $(p_{1},p_{2})$ separates $p$ and $q$. Hence $p,p_{1},p_{2}$ and $q$ are cocircular. Then $\angle p_{1}qp=\angle p_{1}p_{2}p=\pi/3$, so $[pq]$ is a continuation of the remaining edge adjacent to $q$ in $S$. By Ptolemy’s theorem
@@ 2.4 Stability of finite solutions
Let $\mathcal{A}=\{a_{1},\dots,a_{n}\}$, and let the shifted vertices form the set $\mathcal{B}=\{b_{1},\dots,b_{n}\}$, where $b_{i}\in B_{\varepsilon/(5n)}(a_{i})$.
Add segments $a_{i}b_{i}]$ to $\mathcal{S}t$, where $b_{i}$ is the shifted vertex; then replace all edge-segments $a_{i}x]$ with $b_{i}x]$, and consider the unique local minimum $\mathcal{S}t_{new}$ for the terminal set $\mathcal{B}$ from Proposition 1. Each terminal has degree 1 or 2 in $R$, so by the triangle inequality
@@
Consider an arbitrary locally minimal tree $S$ for $\mathcal{B}$ with a topology not in $D(T)$. Similarly to the previous paragraph, add segments $a_{i}b_{i}]$ to $S$, replace all edge-segments $b_{i}x]$ with $a_{i}x]$ and consider the unique local minimum $S_{old}$ for the terminal set $\mathcal{A}$ from Proposition 1. Similarly, each terminal has degree 1, 2 or 3, so
@@
By assumption $\mathcal{H}^{1}(\mathcal{S}t)+\varepsilon<\mathcal{H}^{1}(S_{old})$; putting all together we have the inequality
$\mathcal{H}^{1}(\mathcal{S}t_{new})<\mathcal{H}^{1}(\mathcal{S}t)+2\varepsilon/5<\mathcal{H}^{1}(S_{old})-3\varepsilon/5<\mathcal{H}^{1}(S),$
which finishes the proof. ∎